Jason had saved $60 and plans to save $20 a month, while Jacob had saved $45 and plans to save $40 a month. When will their savings be the same?
Answer by solve_for_x(190)

Let x represent the number of months that Jason and Jacob must save until their amounts are equal.

At that time, Jason's total amount saved will be: 20x + 60

And Jacob's total amount saved will be: 40x + 45

Since the amounts must be the same, set them equal to each other:

40x + 45 = 20x + 60

Subtract 20x from both sides:

40x - 20x + 45 = 60

20x + 45 = 60

Subtract 45 from both sides:

20x + 45 - 45 = 60 - 45

20x = 15

Divide both sides by 20 to get:

x = 15/20

x = 3/4

Their savings will be the same after only 3/4 of a month.